OpenStudy (anonymous):

Use substitution to solve the system -5x+2y=15 -2x+y=7

OpenStudy (anonymous):

First solve for either one of the variables, im going to solve for y from -2x+y=7 we can see that y=7+2x we then plug 7+2x in for y in the first equation, -5x+2(7+2x)=15 We can now solve for x Plug this x value you into either one of the equations and solve for y.

OpenStudy (anonymous):

Thanks. Much Obliged. I gave you a best response!